OpenStudy (anonymous):

Use the elimination method to solve the system of equations. a+6b= -1 a +2b = -5

OpenStudy (anonymous):

subtract one equation from the other. Since there is already the same amount of a's in both equations, this term should drop out and you should be able to solve for b

OpenStudy (anonymous):

8b = -6/8

OpenStudy (anonymous):

Try again

OpenStudy (anonymous):

a + 6b = -1 - (a + 2b = -5) ---------------

OpenStudy (anonymous):

8b=-6

OpenStudy (anonymous):

you are subtracting. The negative sign needs to be distributed to each term

OpenStudy (anonymous):

4b=4

OpenStudy (anonymous):

correct, so what is b?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Correct, so now that we know b, we plug that in to one of the equations to solve for a, it doesn't matter which one. I usually choose the easiest of the 2

OpenStudy (anonymous):

I must note that if 2 equations don't have the same amount of either unknowns, you are allowed to multiply an equation by a scalar to fix this problem. Ex) 2a + 6b = -1 (a + 2b = -5)*2 2a + 6b = -1 2a + 4b = -10 Now you could eliminate. If you wanted to could have multiplied by a -2 so that the equations looked like... 2a + 6b = -1 -2a + -4b = 10 Now you can add the equations instead of subtract. This is what I try to do. I don't like subtraction that much

OpenStudy (anonymous):

a+6(1) = -1 a+6= -1, a =-1 +(-6) a=-7

OpenStudy (anonymous):

correct. good job

OpenStudy (anonymous):

Thanks